# Exploring Community-Powered Conversational Agent for Health Knowledge Acquisition: A Case Study in Colorectal Cancer

### v1

@@ 1. Introduction
With rich information shared and discussed by people with similar experience, online communities have become a popular place for gaining knowledge about health. Based on whether the communities officially involve certified medical professionals, we categorize them into doctor-led communities (e.g., Dingxiang Doctor  [^1] https://dxy.com/ and WebMD [^2] https://www.webmd.com/) and the peer-led communities (e.g., Reddit r/AnalFistula  [^3] https://www.reddit.com/r/AnalFistula/ and RED  [^4] Xiaohongshu in Chinese, https://www.xiaohongshu.com/explore). The doctor-led communities are valuable as users can access professionally validated answers to a wide set of patients’ questions and wiki-like documents that explain the diseases (Ding et al., 2020; Ma et al., 2018; Shlomo Vinker et al., 2007; Nobles et al., 2020), and the peer-led communities are beneficial in terms of informational and emotional support provided by people with similar health challenges (Milton et al., 2023; Yang et al., 2019; Peng et al., 2020; Young and Miller, 2019; Anawade Sr et al., 2024; Nobles et al., 2020). The user groups of these communities include those who are suffering from diseases and actively participate in the online discussion, as well as those who do not necessarily have a disease and just browse the communities to seek health information of their interests (i.e., merely readers). In this paper, we focus on the later user group, especially the younger adults, because compared to other age groups, they are more active users of social media (Chaffey, 2016; Pacheco, 2023; Gil-Clavel and Zagheni, 2019) and more likely to seek and rely on online health information (Fox et al., 2011).
@@
Despite the benefits, people who browse online communities to learn health knowledge often face low-quality and fragmented content (van der Eijk et al., 2013; Nakikj and Mamykina, 2018). Prior HCI researchers have indicated two promising approaches to mitigate this issue. The first approach focuses on data modeling, which involves developing computational models to identify useful content for learning purposes, such as mentioned UI components and visual elements in the comments (Peng et al., 2024) and propositions in the answers about activity plans (Liu et al., 2022). As for the health domain, related work has modeled content in online health communities to understand the members’ interactions, such as the topics and sought social support of the posts (Kim et al., 2023; Wang et al., 2021; De Choudhury and Kiciman, 2017; Chancellor et al., 2018; Peng et al., 2020), as well as the factors that affect support-seekers’ replies to the received comments (Peng et al., 2021; Arguello et al., 2006; Bakhshi et al., 2014). However, little work has targeted browsers who would like to learn health knowledge in online communities and sought to computationally identify the helpful content to support these browsers.
@@
The second approach explores the interaction designs of tools that keep learners’ focuses on high-quality content of their interests. These designs are normally powered by the modeled community data and could include components like conversational agents (CAs) (Peng et al., 2024; Liu et al., 2023), note-taking panels (Liu et al., 2022), mind maps (Chen et al., 2025), and so on. In this paper, we specifically investigate the interaction designs of CAs to support health knowledge acquisition from online communities, because CAs have been demonstrated effective and engaging in various learning tasks (e.g., factual knowledge (Ruan et al., 2019), argumentation writing (Wambsganss et al., 2021), and programming (Winkler et al., 2020; Farah et al., 2024; Ross et al., 2023; Sun et al., 2022; Weisz et al., 2021)). Recent advances in large language models (LLMs) further improve CAs’ capabilities in understanding users’ intentions in their messages and providing structured and relevant responses. However, previous CAs for learning support tasks usually require substantial human effort to prepare specialized knowledge bases and seldom leverage the rich resources from online communities. One of the exceptions is DesignQuizzer (Peng et al., 2024), which prompts multiple-choices questions adapted to comments in an online design community to support learning of visual design. Nevertheless, it leverages a system-questioning design, while users who browse online communities for health knowledge acquisition could be more used to raise questions of their interests, as they usually do when interacting with CAs. Overall, the following research questions (RQs) are under-investigated and guide our work:
@@
We use colorectal cancer, the second most prevalent cancer in China as of 2022 (Han et al., 2024), as a case to implement this workflow in CanAnswer. After data processing with the proposed computational workflow, the data used in CanAnswer includes 24 QA pairs in the based dataset, the 2048 QA pairs in the Disease Lookup and 400 doctor-patient conversations sourced from an doctor-led online community named Dingxiang Doctor, and 246 posts related to colorectal cancer in from a peer-led community named RED. To evaluate CanAnswer, we conducted a between-subjects study with 24 participants who were asked to learn knowledge about colorectal cancer in the lab sessions, using either CanAnswer in the experiment condition or the community interfaces and an LLM in the baseline condition. The results show that participants with CanAnswer recalled significantly more knowledge points after the learning session and were significantly more engaged in the learning process. Participants deemed the suggested community-powered follow-up questions helpful in guiding the learning process. We also interviewed six experts majoring in medicine to gain their feedback on using doctor-led and peer-led communities in CanAnswer. Both the participants in the between-subjects study and experts valued the data from Dingxiang Doctor in providing professional responses and highlighted the practicability and emotional support of real-life examples in RED. However, they express concerns about the credibility of data from RED. We discuss insights of our findings for leveraging community data and designing conversational agents to support health knowledge acquisition.
@@ 2.1. Seeking Health Knowledge in Online Communities
The Internet offers accessible resources for people who seek to self-diagnose and manage their health conditions (White and Horvitz, 2009; Berland et al., 2001; Lee et al., 2014; Eysenbach et al., 2002; Shahsavar et al., 2023; Becker, 2004). Among various online channels (e.g., government health agencies, video platforms), online communities are unique because they bring together a large scale of members with similar interests to share or exchange information with each other (Milton et al., 2023; Ernala et al., 2018; Young and Miller, 2019). We categorized online communities into doctor-led, where licensed professionals create or moderate content, and peer-led, where lay users share personal experiences and support each other.
@@
Doctor-led online communities play a key role in delivering accurate, reliable medical information. These platforms often provide expert-reviewed articles, treatment guidelines, and Q&A services with certified doctors. Ding et al. (2020) conducted an empirical study on Ask the Doctor (AtD), an online health platform in China facilitating patient-provider communication. They found AtD to be a valuable supplement to offline healthcare and highlighted learning and education as promising directions (Ding et al., 2020). Building on these insights, recent research has explored advanced applications in doctor-led communities, such as AI-powered Clinical Decision Support (CDS) systems that offer personalized treatment recommendations for diabetes based on large-scale patient data (Burgess et al., 2023). These innovations underscore the potential of professionally curated communities to integrate technology into health management.
@@
Peer-led communities, in contrast, are primarily characterized by the exchange of personal experiences and the provision of emotional support—such as kindness, encouragement, and empathy—among members facing shared health challenges (Sharma and De Choudhury, 2018). Although the information shared may lack professional verification, it offers experiential insights that are especially valuable for managing chronic conditions and addressing emotional needs through peer support (Sinha et al., 2018; Kendal et al., 2017; Rains and Young, 2009; Milton et al., 2023). Studies on such communities have shown that perceived similarity among peers can foster trust and improve health outcomes by enabling more relatable and actionable advice (Frost et al., 2008). Long-term engagement in such communities has been associated with reduced depression, enhanced self-efficacy, and improved quality of life (Farnood et al., 2022; Attard and Coulson, 2012). This dual approach—combining professional guidance with peer support—has been shown to significantly enhance users’ ability to manage their health effectively.
@@
Despite the benefits of these communities for learning about diseases, users can encounter low-quality and irrelevant content. As noted by Kanthawala et al. (2016), health-related content in online communities is predominantly based on personal experiences rather than clinical guidelines, which may not always conform to established clinical standards. This concern is echoed in Farnood et al. (2020)’s review of 25 studies that assess perceptions of online self-diagnosis and health information seeking. The review found that, while participation in online cancer communities can positively impact patient-reported outcomes, the quality of discussions and shared information varied widely, influenced by factors such as the type of information sought, the source, and the levels of clinical accuracy and relevance within community discussions (Farnood et al., 2020). Our work is motivated by the health knowledge and support users seek in both doctor-led and peer-led online communities. We propose a computational workflow to help users effectively access high-quality, relevant information and emotional support shared in these spaces.
@@ 2.2. Computational Methods for Modeling Community Data
Existing literature has leveraged computational methods to model community data, e.g., posts and comments, for various purposes in different domains. For example, Erickson and Yan (2024) used the Natural Language Toolkit to tokenize the Facebook posts and Instagram pages to understand how the introduction of Facebook Reactions influenced the behavior of the 114th US Congress. Liu et al. (2023) designed a processing pipeline for extracting and summarizing augmentative elements from question threads to power a tool that helps lurkers contribute to the question-answering community. In the health domains, prior HCI work primarily focuses on understanding and facilitating user interaction in the communities (Peng et al., 2021; Sharma and De Choudhury, 2018; Yang et al., 2019; Jin et al., 2023). For instance, Yang et al. (2019) built an unsupervised Gaussian mixture model from the data to discover 11 roles members occupy in an online support community for cancer patients. This model supports the analyses of how roles change over members’ lifecycles and how roles predict long-term participation in the community. Peng et al. (2020) built random forest and XGBoost models to assess the amount of informational and emotional support conveyed in the comments, which helps people draft more supportive remarks to others in an online depression community. To leverage community data for learning purposes, one closely related work conducted by Peng et al. (2024) proposed a computational workflow that extracts critiques, suggestions, and rationales on visual elements (e.g., color, layout) from comments in a visual design community. Nevertheless, little work has targeted browsers who aim to acquire health-related knowledge in online communities, nor has it explored how to computationally model community data for learning about diseases, which, compared to the visual design learning tasks, could be more sensitive to the information quality, source credibility, and emotional support of the community content (Jin et al., 2016). In this paper, we collaborate with a team of doctors to develop a computational workflow that processes community data for health-related learning tasks. We also conduct user studies to understand how users perceive the online communities for learning about diseases.
@@ 2.3. Conversational Agents in Education and Health Domains
Previous research has developed a set of conversational agents (CAs) to engage users in learning tasks. For instance, Wambsganss et al. (2021) developed ArgueTutor that judges the argumentative writing performance of users’ essays and suggests how to improve. Winkler et al. (2020) developed Sara, which acts like a teacher by asking students questions about programming during an online video lecture. However, little work explores the usage of community data to power a CA for learning tasks. One closely related CA is DesignQuizzer (Peng et al., 2024), which prompts multiple-choices questions about user interface design examples curated from online design communities. Our proposed CanAnswer is inspired by these CAs but takes the role of an question responder instead of a quizzer, as users often raise questions to seek information in online communities (Ding et al., 2020).
Apart from the fields of learning and education, CAs have played various roles of social actors in health domains. For example, Li et al. (2024a) proposed a pre-consultation conversational agent that combines questions from existing pre-consultation questionnaires and asks participants questions similar to those in physician-initiated text-based consultations. Furthermore, Park et al. (2024) developed a CA to help children with ADHD improve their executive function. The CA incorporates sound alarms, voice interactions, and motivational mechanisms such as awarding star stickers, along with sound and visual cues to guide users in completing tasks. Ding et al. (2022) developed TalkTive, a chatbot using backchanneling techniques to engage older adults in conversations during neurocognitive disorder screenings for further professional assessments. However, most health CAs focus solely on assisting users in completing specific medical tasks, such as pre-consultation (Li et al., 2024a; Te Pas et al., 2020; Ni et al., 2017; Li, 2024), improving executive function for ADHD (Park et al., 2024), and screening for Alzheimer’s disease (Talktive (Ding et al., 2022)). Besides, these CAs usually rely on substantial effort from doctors to prepare the database. Differently, our CanAnswer primarily focuses on popularizing health knowledge based on a large amount of community data and a small amount of doctors’ input.
@@ 3. Formative Study
To address RQ1, we conducted a formative study to investigate users’ challenges and expected features when acquiring health content in online communities. Inspired by (Yuan et al., 2023), we first brainstormed potential features based on existing literature, and then derived design goals for a CA by incorporating findings from the formative study.
@@ 3.1. Participants
We recruited 56 participants (S1–56, 21 Males, 35 Females) through advertisements posted in an online community at a Chinese university and via word of mouth. All participants were native Chinese speakers, and the study was conducted entirely in Chinese. Most participants were students, including 37 undergraduates, 18 master’s students and one non-student participant. Participants ranged in age from 18 to 51 years (M = 22.32, SD = 4.34), representing our primary target user group, as young individuals increasingly need to acquire health knowledge for early prevention and prognosis (Nagai and Kim, 2017). Our participants came from a range of academic disciplines, including science, engineering, medicine, humanities, and business. Table 1 presents the demographic information of six participants (P1-6, 3 Males, 3 Females) who took part in our interview study, whose ages range from 22 to 51 (Mean = 27.2, SD = 10.7). These participants were recruited via follow-up invitations from the survey study who had expressed interest in continued participation. P2, P3, and P4 had experiences searching for health information in online communities because their immediate relatives (e.g., mother and uncle) suffered or were suffering from cancers, including colorectal cancer. The other three participants used to search for health information online as they felt sick, majored in medicine or were interested in learning about it.

### v2

@@ 1. Introduction
With rich information shared and discussed by people with similar experience, online communities have become a popular place for gaining knowledge about health. Based on whether the communities officially involve certified medical professionals, we categorize them into doctor-led communities (e.g., Dingxiang Doctor  [^1] https://dxy.com/ and WebMD [^2] https://www.webmd.com/) and the peer-led communities (e.g., Reddit r/AnalFistula  [^3] https://www.reddit.com/r/AnalFistula/ and RED  [^4] Xiaohongshu in Chinese, https://www.xiaohongshu.com/explore). The doctor-led communities are valuable as users can access professionally validated answers to a wide set of patients’ questions and wiki-like documents that explain the diseases (ask_the_doctor; ma2018professional; Vinker01012007; nobles2020examining), and the peer-led communities are beneficial in terms of informational and emotional support provided by people with similar health challenges (similar_situation; emotional_informational1; emotional_informational2; rare_disease; anawade2024connecting; nobles2020examining). The user groups of these communities include those who are suffering from diseases and actively participate in the online discussion, as well as those who do not necessarily have a disease and just browse the communities to seek health information of their interests (i.e., merely readers). In this paper, we focus on the later user group, especially the younger adults, because compared to other age groups, they are more active users of social media (chaffey2016global; social_media; social_media2) and more likely to seek and rely on online health information (fox2011social).
@@
Despite the benefits, people who browse online communities to learn health knowledge often face low-quality and fragmented content (fragmented; overload_community). Prior HCI researchers have indicated two promising approaches to mitigate this issue. The first approach focuses on data modeling, which involves developing computational models to identify useful content for learning purposes, such as mentioned UI components and visual elements in the comments (DesignQuizzer) and propositions in the answers about activity plans (Planhelper). As for the health domain, related work has modeled content in online health communities to understand the members’ interactions, such as the topics and sought social support of the posts (kim2023supporters; wang2021cass; de2017language; chancellor2018norms; emotional_informational2), as well as the factors that affect support-seekers’ replies to the received comments (emotional_informational; arguello2006talk; bakhshi2014faces). However, little work has targeted browsers who would like to learn health knowledge in online communities and sought to computationally identify the helpful content to support these browsers.
@@
The second approach explores the interaction designs of tools that keep learners’ focuses on high-quality content of their interests. These designs are normally powered by the modeled community data and could include components like conversational agents (CAs) (DesignQuizzer; liu2023coargue), note-taking panels (Planhelper), mind maps (chen2025redesignonlinedesigncommunities), and so on. In this paper, we specifically investigate the interaction designs of CAs to support health knowledge acquisition from online communities, because CAs have been demonstrated effective and engaging in various learning tasks (e.g., factual knowledge (ruan2019quizbot), argumentation writing (wambsganss2021arguetutor), and programming (programming_bot; programming_bot2; programming_bot3; programming_bot4; programming_bot5)). Recent advances in large language models (LLMs) further improve CAs’ capabilities in understanding users’ intentions in their messages and providing structured and relevant responses. However, previous CAs for learning support tasks usually require substantial human effort to prepare specialized knowledge bases and seldom leverage the rich resources from online communities. One of the exceptions is DesignQuizzer (DesignQuizzer), which prompts multiple-choices questions adapted to comments in an online design community to support learning of visual design. Nevertheless, it leverages a system-questioning design, while users who browse online communities for health knowledge acquisition could be more used to raise questions of their interests, as they usually do when interacting with CAs. Overall, the following research questions (RQs) are under-investigated and guide our work:
@@
We use colorectal cancer, the second most prevalent cancer in China as of 2022 (cancer_rank), as a case to implement this workflow in CanAnswer. After data processing with the proposed computational workflow, the data used in CanAnswer includes 24 QA pairs in the based dataset, the 2048 QA pairs in the Disease Lookup and 400 doctor-patient conversations sourced from an doctor-led online community named Dingxiang Doctor, and 246 posts related to colorectal cancer in from a peer-led community named RED. To evaluate CanAnswer, we conducted a between-subjects study with 24 participants who were asked to learn knowledge about colorectal cancer in the lab sessions, using either CanAnswer in the experiment condition or the community interfaces and an LLM in the baseline condition. The results show that participants with CanAnswer recalled significantly more knowledge points after the learning session and were significantly more engaged in the learning process. Participants deemed the suggested community-powered follow-up questions helpful in guiding the learning process. We also interviewed six experts majoring in medicine to gain their feedback on using doctor-led and peer-led communities in CanAnswer. Both the participants in the between-subjects study and experts valued the data from Dingxiang Doctor in providing professional responses and highlighted the practicability and emotional support of real-life examples in RED. However, they express concerns about the credibility of data from RED. We discuss insights of our findings for leveraging community data and designing conversational agents to support health knowledge acquisition.
@@ 2.1. Seeking Health Knowledge in Online Communities
The Internet offers accessible resources for people who seek to self-diagnose and manage their health conditions (white2009experiences; berland2001health; lee2014dr; eysenbach2002empirical; shahsavar2023user; old_adult_resource). Among various online channels (e.g., government health agencies, video platforms), online communities are unique because they bring together a large scale of members with similar interests to share or exchange information with each other (similar_situation; similar_situation2; rare_disease). We categorized online communities into doctor-led, where licensed professionals create or moderate content, and peer-led, where lay users share personal experiences and support each other.
@@
Doctor-led online communities play a key role in delivering accurate, reliable medical information. These platforms often provide expert-reviewed articles, treatment guidelines, and Q&A services with certified doctors. ask_the_doctor conducted an empirical study on Ask the Doctor (AtD), an online health platform in China facilitating patient-provider communication. They found AtD to be a valuable supplement to offline healthcare and highlighted learning and education as promising directions (ask_the_doctor). Building on these insights, recent research has explored advanced applications in doctor-led communities, such as AI-powered Clinical Decision Support (CDS) systems that offer personalized treatment recommendations for diabetes based on large-scale patient data (burgess2023healthcare). These innovations underscore the potential of professionally curated communities to integrate technology into health management.
@@
Peer-led communities, in contrast, are primarily characterized by the exchange of personal experiences and the provision of emotional support—such as kindness, encouragement, and empathy—among members facing shared health challenges (emotional_support). Although the information shared may lack professional verification, it offers experiential insights that are especially valuable for managing chronic conditions and addressing emotional needs through peer support (sinha2018use; kendal2017moderated; meta-analysis; similar_situation). Studies on such communities have shown that perceived similarity among peers can foster trust and improve health outcomes by enabling more relatable and actionable advice (frost2008social). Long-term engagement in such communities has been associated with reduced depression, enhanced self-efficacy, and improved quality of life (farnood2022understanding; attard2012thematic). This dual approach—combining professional guidance with peer support—has been shown to significantly enhance users’ ability to manage their health effectively.
@@
Despite the benefits of these communities for learning about diseases, users can encounter low-quality and irrelevant content. As noted by kanthawala2016answers, health-related content in online communities is predominantly based on personal experiences rather than clinical guidelines, which may not always conform to established clinical standards. This concern is echoed in farnood2020mixed’s review of 25 studies that assess perceptions of online self-diagnosis and health information seeking. The review found that, while participation in online cancer communities can positively impact patient-reported outcomes, the quality of discussions and shared information varied widely, influenced by factors such as the type of information sought, the source, and the levels of clinical accuracy and relevance within community discussions (farnood2020mixed). Our work is motivated by the health knowledge and support users seek in both doctor-led and peer-led online communities. We propose a computational workflow to help users effectively access high-quality, relevant information and emotional support shared in these spaces.
@@ 2.2. Computational Methods for Modeling Community Data
Existing literature has leveraged computational methods to model community data, e.g., posts and comments, for various purposes in different domains. For example, erickson2024affective used the Natural Language Toolkit to tokenize the Facebook posts and Instagram pages to understand how the introduction of Facebook Reactions influenced the behavior of the 114th US Congress. liu2023coargue designed a processing pipeline for extracting and summarizing augmentative elements from question threads to power a tool that helps lurkers contribute to the question-answering community. In the health domains, prior HCI work primarily focuses on understanding and facilitating user interaction in the communities (emotional_informational; emotional_support; emotional_informational1; jin2023understanding). For instance, emotional_informational1 built an unsupervised Gaussian mixture model from the data to discover 11 roles members occupy in an online support community for cancer patients. This model supports the analyses of how roles change over members’ lifecycles and how roles predict long-term participation in the community. emotional_informational2 built random forest and XGBoost models to assess the amount of informational and emotional support conveyed in the comments, which helps people draft more supportive remarks to others in an online depression community. To leverage community data for learning purposes, one closely related work conducted by DesignQuizzer proposed a computational workflow that extracts critiques, suggestions, and rationales on visual elements (e.g., color, layout) from comments in a visual design community. Nevertheless, little work has targeted browsers who aim to acquire health-related knowledge in online communities, nor has it explored how to computationally model community data for learning about diseases, which, compared to the visual design learning tasks, could be more sensitive to the information quality, source credibility, and emotional support of the community content (adoption_decision). In this paper, we collaborate with a team of doctors to develop a computational workflow that processes community data for health-related learning tasks. We also conduct user studies to understand how users perceive the online communities for learning about diseases.
@@ 2.3. Conversational Agents in Education and Health Domains
Previous research has developed a set of conversational agents (CAs) to engage users in learning tasks. For instance, wambsganss2021arguetutor developed ArgueTutor that judges the argumentative writing performance of users’ essays and suggests how to improve. programming_bot developed Sara, which acts like a teacher by asking students questions about programming during an online video lecture. However, little work explores the usage of community data to power a CA for learning tasks. One closely related CA is DesignQuizzer (DesignQuizzer), which prompts multiple-choices questions about user interface design examples curated from online design communities. Our proposed CanAnswer is inspired by these CAs but takes the role of an question responder instead of a quizzer, as users often raise questions to seek information in online communities (ask_the_doctor).
Apart from the fields of learning and education, CAs have played various roles of social actors in health domains. For example, pre-consultation proposed a pre-consultation conversational agent that combines questions from existing pre-consultation questionnaires and asks participants questions similar to those in physician-initiated text-based consultations. Furthermore, ADHD_child developed a CA to help children with ADHD improve their executive function. The CA incorporates sound alarms, voice interactions, and motivational mechanisms such as awarding star stickers, along with sound and visual cues to guide users in completing tasks. ding2022talktive developed TalkTive, a chatbot using backchanneling techniques to engage older adults in conversations during neurocognitive disorder screenings for further professional assessments. However, most health CAs focus solely on assisting users in completing specific medical tasks, such as pre-consultation (pre-consultation; pre-consultation_2; pre-consultation_3; pre_consultation4), improving executive function for ADHD (ADHD_child), and screening for Alzheimer’s disease (Talktive (ding2022talktive)). Besides, these CAs usually rely on substantial effort from doctors to prepare the database. Differently, our CanAnswer primarily focuses on popularizing health knowledge based on a large amount of community data and a small amount of doctors’ input.
@@ 3. Formative Study
To address RQ1, we conducted a formative study to investigate users’ challenges and expected features when acquiring health content in online communities. Inspired by (yuan2023critrainer), we first brainstormed potential features based on existing literature, and then derived design goals for a CA by incorporating findings from the formative study.
@@ 3.1. Participants
We recruited 56 participants (S1–56, 21 Males, 35 Females) through advertisements posted in an online community at a Chinese university and via word of mouth. All participants were native Chinese speakers, and the study was conducted entirely in Chinese. Most participants were students, including 37 undergraduates, 18 master’s students and one non-student participant. Participants ranged in age from 18 to 51 years (M = 22.32, SD = 4.34), representing our primary target user group, as young individuals increasingly need to acquire health knowledge for early prevention and prognosis (cancer_prevent_early_age). Our participants came from a range of academic disciplines, including science, engineering, medicine, humanities, and business. Table 1 presents the demographic information of six participants (P1-6, 3 Males, 3 Females) who took part in our interview study, whose ages range from 22 to 51 (Mean = 27.2, SD = 10.7). These participants were recruited via follow-up invitations from the survey study who had expressed interest in continued participation. P2, P3, and P4 had experiences searching for health information in online communities because their immediate relatives (e.g., mother and uncle) suffered or were suffering from cancers, including colorectal cancer. The other three participants used to search for health information online as they felt sick, majored in medicine or were interested in learning about it.